Main INDEX
Monthly INDEX
PREV
NEXT
User name LeRose
Log entry time 16:11:29 on June23,2005
Entry number 147015
Followups:
keyword=Shift Summary
Thursday June 23, 2005 Day
J. LeRose
D. Whitbeck
At start there are problems with the energy spread of the beam (as reported by the SLI).
Beam will go away at 08:30 and probably be gone until ~13:30
08:44 beam off for beam studies. We go to restricted access
Techs replace contoller on Q1 right (suspected cause of noisy leadflow readback that was giving false alarms)
Sirish works on Compton Laser comes out at 12:10 and declares it all works "problem solved"
Nilanga and student do some work on their detectors
Riad works on his Lumi stuff.
12:15 Left Dipole tripped and right septum
13:45 MCC is having problems repairing an arc power supply. Will be down for 30-60 minutes longer. Techs checking on right Q1 lead flow again (earlier fix appears not to have worked)
14:25 Left arm controls froze. Rebooted IOCha14 (see halog 147011)
14:20 Eugene and Roman in the Hall for Moller work
15:27 MCC ready to start delivery. Pulled Eugene and Roman out of the Hall.
16:00 Waiting!
run start/stop events
4842 08:10/08:44 0.75M stopped for beam studies, energy spread not good